the number of four digit even numbers that can be formed using 0,1,2,3,4,5,6 without repetition is what.

Grade:12

2 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
For an even number its unit digit should be 0,2,4 or 6 so there are 4 ways.
For a 4 digit number __ __ __ __
6 5 4 4
As it is a 4 digit number 1stnumber shouldnot be zero.so 6 ways
Hence total number of ways=6*5*4*4=480 ways
Thanks and Regards
Shaik Aasif
askIITians faculty
Latika Leekha
9 years ago
There can be two cases:
Case (1): When the first digit is even:
We can have the first digit as 2 or 4 or 6 = 3 choices
Similarly, the last digit can be any of 0, 2, 4 or 6 except the first chosen = 3 choices.
Hence, total number of possibilities = 3. 5. 4.3 = 180
Case (2): When the first digit is odd:
We can have the first digit as 1, 3 or 5 = 3 choices
The last digit can be = 0, 2, 4, 6 = 4 choices
Hence, the total number of possibilities = 3.5.4.4. = 240
Hence, the total number of cases = 180 + 240 = 420 ways.
Thanks & Regards
Latika Leekha
askIITians Faculty